What is the concentration of LiCl (lithium chloride) when 2.5g of LiCl are dissolved in water and the volume of the final solution is 45.0mL? The formula for %(m/v) is shown below.
Answer #1

Answer:

Given,

To determine the concentration of Licl

weight of the lithium chloride = 2.5g

Molar mass of the lithium chloride = 42.394g/mol

Now,

Mole of the Lithium chloride = mass/volume

Moles of Licl = 2.5/42.394= 0.058971

Now consider,

Concentration = mole/volume in litre

Concentration = 0.058971/(45*10^-3)

Licl concentration= 1.3105mol/litre